# SIC-Free Rate-Splitting Multiple Access: Constellation-Constrained Optimization and Application to Large-Scale Systems

### v1

@@ VI-C Weighted Sum-Rate Performance
(a) $N_{T}=2$, $K=2$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}}\}=\{0,\pi/9\}$, $\kappa=10\text{ dB}$.
(b) $N_{T}=4$, $K=3$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}},\theta_{3}^{\text{az}}\}=\{-\pi/18,0,\pi/18\}$, $\kappa=10\text{ dB}$.
Figure 3: Ergodic sum-rate performance.
We first evaluate the sum-rate performance by setting $[u_{1},...,u_{k}]=\mathbf{1}^{1\times K}$. Figure 3 depicts the ergodic sum-rate of SIC-based and SIC-free RSMA under finite constellations. The performance of SDMA under finite constellations is also depicted by fixing the transmission mode to be the first one in each dictionaries. Furthermore, the performance of RSMA and SDMA under Gaussian distributed input signals, generated following [6], is included as a reference. It can be observed that, in all cases, SIC-free RSMA preserves most of the superiority of SIC-based RSMA over SDMA, which is consistent with the conclusion in [19].
@@
Figure 4 depicts the ratio of the average common stream power to the transmit power budget with SIC-based and SIC-free RSMA, i.e. $\mathbb{E}\left[\|\mathbf{p}_{\text{c}}\|^{2}/P_{\text{T}}\right]$, under the same setup as in Figure 3 (a). This result visualizes the fact that, despite SIC-based and SIC-free achieving similar performance, they demand very different precoder designs and power allocation strategies. Notably, SIC-free RSMA tends to allocate less power to the common stream in low SNR regime. This is to compensate for the fact that non-SIC receivers are not able to remove interference caused by the common stream. It can be observed that the algorithm tends to turn off the common stream at extremely high and low SNRs. It can also be observed that as $R_{\text{max}}$ increases, the power allocation curves shift to the right-hand side, because a higher throughput needs to be supported by a higher SNR.
@@
By applying various weights to the users, the rate region can be obtained. Figure 6 depicts the ergodic rate region with $N_{\text{T}}=2$ and $K=2$ and indicates that RSMA achieves better rate regions than SDMA under the given finite-constellation setup, similar to Gaussian distributed inputs [6]. It also suggests that SIC-free RSMA performs similarly to SIC-based RSMA under different weights; hence, the two achieve similar rate regions. This result extends the conclusion in [19] from sum-rate to rate regions.
(a) $N_{\text{T}}=2$, $K=2$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}}\}=\{0,\pi/9\}$, $\kappa=10\text{ dB}$.
(b) $N_{\text{T}}=4$, $K=3$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}},\theta_{3}^{\text{az}}\}=\{-\pi/18,0,\pi/18\}$, $\kappa=10\text{ dB}$.
@@ VI-D Max-Min Fairness Performance
Figure 7 depicts the MMF performance of SIC-based and SIC-free RSMA under finite constellations. The performance of SDMA under finite constellations is also depicted by fixing the transmission mode as in the WSR evaluations. In addition, the MMF performance of RSMA and SDMA under Gaussian distributed input signals, following the work in [42], is included as a reference. It can be observed that RSMA schemes achieve better MMF performance than SDMA under finite constellations; meanwhile, SIC-free RSMA preserves most of the MMF performance of SIC-based RSMA. This result extends the conclusion in [19] from sum-rate to MMF evaluations.
@@ VI-E Performance in Large-Scale Systems
We consider a large-scale rectangular antenna array with $N_{\text{T}}=128$, with 16 antennas in each row along the $y$-axis and 8 antennas in each column along the $z$-axis, serving 64 users. Figures 8 and 9 depict the sum-rate and MMF performance of SDMA, SIC-based and SIC-free RSMA for large-scale systems, which is enabled by techniques including user grouping, null space projection, and dimensionality reduction introduced in Section V. For the channel model, $\theta_{k}^{\text{az}}$ and $\theta_{k}^{\text{el}}$, $\forall k\in\mathcal{K}$, are randomly drawn from the range of $[-\pi/4,\pi/4]$ and $[-\pi/6,\pi/6]$ with uniform distribution. $\kappa$ is set to 20dB to reflect the sparse nature of channels in typical large-scale antenna arrays. It can be observed that the conclusions drawn from small-scale systems as discussed previously, i.e. RSMA schemes outperform SDMA in both sum-rate and MMF, and SIC-free RSMA preserves most of the superiority of SIC-based RSMA, still holds in large-scale systems.
@@ VII Conclusion
In conclusion, in this work, we aimed to investigate the fundamental limits of SIC-based and SIC-free RSMA under finite constellations. We first derived the rate expressions for SIC-based and SIC-free RSMA. We then proposed precoding optimization algorithms for WSR and MMF optimization problems. To enable these optimization algorithms in large-scale systems with feasible computational burden, we introduced user grouping, null space projection and dimensionality reduction techniques. Numerical results show that RSMA schemes outperform SDMA under finite constellations with either WSR or MMF optimization, similar to the Gaussian distributed inputs. Furthermore, SIC-free RSMA preserves most of the superiority of SIC-based RSMA, which extends the conclusion in [19] from low-complexity precoders to fully optimized precoders, from sum-rate objective to WSR and MMF objectives, and from small-scale systems to large-scale systems.
@@ Appendix A Proof of Proposition 1
Since $\mathbf{P}$ is given and the private rates, $R_{\text{p},k},\forall k\in\mathcal{K}$, are not functions of $\mathbf{c}$, $\mathcal{P}_{1}$ is equivalent to
$\begin{split}\underset{\mathbf{c}}{\text{max}}\>\>&\sum_{k\in\mathcal{K}}u_{k}C_{k}\\ \text{s.t.}\>\>\>\;&\sum_{k\in\mathcal{K}}C_{k}\leq R_{\text{c}}\\ &\mathbf{c}\geq\mathbf{0},\end{split}$ (38)
@@ Appendix B Proof of Proposition 2
$\mathcal{P}_{4}$ can be written as follows
$\begin{split}\mathcal{P}_{7}:\;\;\max_{\mathbf{c}}&\;\;\min_{k}C_{k}+R_{\text{p},k}\\ \text{s.t.}&\;\;\;\;\mathbf{c}\succcurlyeq\mathbf{0},\\ &\;\;\;\;\mathbf{1}^{T}\mathbf{c}=R_{\text{c}}.\\ \end{split}$ (39)
We introducing a slack variable, $\eta$, and convert $\mathcal{P}_{7}$ into
$\begin{split}\mathcal{P}_{8}:\;\;\max_{\mathbf{c},t}&\;\;\;\;\;\;\eta\\ \text{s.t.}&\;\;\;\;\mathbf{c}\succcurlyeq\mathbf{0},\\ &\;\;\;\;\mathbf{1}^{T}\mathbf{c}=R_{\text{c}}.\\ &\;\;\;\;C_{k}+R_{\text{p},k}\geq\eta,\forall k.\end{split}$ (40)
Let $(\mathbf{c}^{\star},\eta^{\star})$ denote the solution of $\mathcal{P}_{8}$. Without loss of generality, we assume $R_{\text{p},1}<R_{\text{p},2}<...<R_{\text{p},K}$, and the solution of $\mathcal{P}_{8}$ leads to $C_{k}^{\star}>0$ for $k=1,...,k^{\prime}$ with $1\leq k^{\prime}\leq K$, and $C_{k}^{\star}=0$ for $k=k^{\prime}+1,...,K$. We have the following lemma.
@@
$C_{k}^{\star}+R_{\text{p},k}=\eta^{\star}$ for $k=1,...,k^{\prime}$ and some value for $\eta^{\star}$.
@@
It is trivial to see that $R_{\text{p},k^{\prime}}<\eta^{\star}$. Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be a partition of $\{1,...,k^{\prime}\}$. Suppose $(\mathbf{c}^{\star},\eta^{\star})$ leads to $C^{\star}_{k}+R_{\text{p},k}>\eta^{\star}$ for $k\in\mathcal{K}_{1}$, and $C^{\star}_{k}+R_{\text{p},k}=\eta^{\star}$ for $k\in\mathcal{K}_{2}$. Then, another choice of $\mathbf{c}$, denoted by $\mathbf{c}^{\diamond}$, can be found by $\mathbf{c}^{\diamond}=\alpha\mathbf{c^{\circ}}$, where $C^{\circ}_{k}=(C^{\star}_{k}-R_{\text{p},k}+\eta^{\star})/2$ for $k\in\mathcal{K}_{1}$, $C^{\circ}_{k}=C_{k}^{\star}$ for $k\in\mathcal{K}_{2}$, and $\alpha=R_{\text{c}}/\mathbf{1}^{T}\mathbf{c}^{\circ}$. It is trivial to see that $\alpha>1$. Therefore, $\min_{k\in\{1,...,k^{\prime}\}}C^{\diamond}_{k}+R_{\text{p},k}>\min_{k\in\{1,...,k^{\prime}\}}C^{\star}_{k}+R_{\text{p},k}$, which indicates that $\mathbf{c}^{\star}$ is not the solution of $\mathcal{P}_{11}$. By contradiction, Lemma 2 is proved. ∎
@@
$\mathbf{c}^{\prime}+\mathbf{d}=\eta\mathbf{1},$ (41)
where $\mathbf{d}=[R_{\text{p},1},...,R_{\text{p},k^{\prime}}]^{T}$, and $\mathbf{c}^{\prime}=[C_{1},...,C_{k^{\prime}}]$. Combining (41) and the constraint on $\mathbf{1}^{T}\mathbf{c}=R_{\text{c}}$, the global optimal solution of $\mathbf{c}^{\prime}$ and $\eta$, denoted by $\mathbf{c}^{\prime\star}$ and $\eta^{\star}$, can be then obtained by
$\mathbf{c}^{\prime\star}=\eta^{\star}\mathbf{1}-\mathbf{d},$ (42)
@@
$\eta^{\star}=\frac{R_{\text{c}}+\mathbf{1}^{T}\mathbf{d}}{k^{\prime}}.$ (43)
Although $k^{\prime}$, i.e. the number of users whose are strictly served by the common stream, is still not known, we can test it using (42) and (43). This is because allocating the common stream rate based on (42) and (43) to more users than $k^{\prime}$ will lead to $\mathbf{c}^{\prime}$ containing negative entries, which is infeasible. Hence, by iteratively guessing the number of users to be served by the common stream, Alg. 3 is obtained.
@@ Appendix C Transmission Mode Dictionaries Used For Numerical Results
The transmission mode dictionaries used to produce the numerical results in Section VI are depicted as follows.

### v2

@@ VI-C Weighted Sum-Rate Performance
(a) $N_{T}=2$, $K=2$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}}\}=\{0,\pi/18\}$, $\kappa=10\text{ dB}$.
(b) $N_{T}=4$, $K=3$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}},\theta_{3}^{\text{az}}\}=\{-\pi/36,0,\pi/36\}$, $\kappa=10\text{ dB}$.
Figure 3: Ergodic SR performance.
We first evaluate the sum-rate performance by setting $[u_{1},...,u_{k}]=\mathbf{1}^{1\times K}$. Figure 3 depicts the ergodic sum-rate of SIC-based and SIC-free RSMA under finite constellations. The performance of SDMA under finite constellations is also depicted by fixing the transmission mode to be the first one in each dictionaries. For two users, we also included power-domain Non-Orthogonal Multiple Access (NOMA) [44] as a benchmark, where both users apply the same constellation as in SDMA, and the user ordering is determined by the channel strength [^9] For example, if $\|\mathbf{h}_{1}\|^{2}\geq\|\mathbf{h}_{2}\|^{2}$, user-$1$ decodes user-$2$’s message, applies SIC and decodes its own message, whereas user-$2$ decodes its own message directly.. Furthermore, the performance of RSMA and SDMA under Gaussian distributed input signals, following [6], is included as a reference. It can be observed that, in all cases, SIC-free RSMA preserves most of the superiority of SIC-based RSMA over SDMA and NOMA, which is consistent with the conclusion in [19].
@@
Figure 4 depicts the ratio of the average common stream power to the transmit power budget with SIC-based and SIC-free RSMA, i.e., $\mathbb{E}\left[\|\mathbf{p}_{\text{c}}\|^{2}/P_{\text{T}}\right]$, under the same setup as in Figure 3 (a). This result visualizes the fact that, despite SIC-based and SIC-free achieving similar performance, they demand very different precoder designs and power allocation strategies. Notably, SIC-free RSMA tends to allocate less power to the common stream in low SNR regime. This is to compensate for the fact that SIC-free receivers are not able to remove interference caused by the common stream. It can be observed that the algorithm tends to turn off the common stream at extremely high and low SNRs. It can also be observed that as $R_{\text{max}}$ increases, the power allocation curves shift to the right-hand side, because a higher achievable rate needs to be supported by a higher SNR.
@@
By applying various weights to the users, the rate region can be obtained. Figure 6 depicts the ergodic rate region with $N_{\text{T}}=2$ and $K=2$ and indicates that RSMA achieves better rate regions than SDMA under finite constellations, similar to Gaussian distributed inputs [6]. It also suggests that SIC-free RSMA performs similarly to SIC-based RSMA under different weights; hence, the two achieve similar rate regions. This result extends the conclusion in [19] from sum-rate to rate regions.
(a) $N_{\text{T}}=2$, $K=2$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}}\}=\{0,\pi/18\}$, $\kappa=10\text{ dB}$.
(b) $N_{\text{T}}=4$, $K=3$, $\{\theta_{1}^{\text{az}},\theta_{2}^{\text{az}},\theta_{3}^{\text{az}}\}=\{-\pi/36,0,\pi/36\}$, $\kappa=10\text{ dB}$.
@@ VI-D Max-Min Fairness Performance
Figure 7 depicts the MMF performance of SIC-based and SIC-free RSMA under finite constellations. The performance of SDMA under finite constellations is also depicted by fixing the transmission mode as in the WSR evaluations. For two users, we also included power-domain NOMA as a benchmark, where both users apply the same constellation as in SDMA, and the user ordering is determined by the channel strengthfootnotemark: . In addition, the MMF performance of RSMA and SDMA under Gaussian distributed input signals, following the work in [45], is included as a reference. It can be observed that RSMA schemes achieve better MMF performance than SDMA and NOMA under finite constellations; meanwhile, SIC-free RSMA preserves most of the MMF performance of SIC-based RSMA. This result extends the conclusion in [19] from sum-rate to MMF evaluations.
@@ VI-E Performance in Large-Scale Systems
We consider a large-scale rectangular antenna array with $N_{\text{T}}=128$, with 16 antennas in each row along the $y$-axis and 8 antennas in each column along the $z$-axis, serving 64 users. Figures 8 and 9 depict the sum-rate and MMF performance of SDMA, SIC-based and SIC-free RSMA for large-scale systems, which is enabled by techniques including user grouping, null space projection, and dimensionality reduction introduced in Section V. We also evaluate the impact of user grouping by comparing the performance under the proposed grouping method in Algorithm 5 and random grouping, respectively labeled as ”(ordered)” and ”(unordered)” in the legends. For the channel model, $\theta_{k}^{\text{az}}$ and $\theta_{k}^{\text{el}}$, $\forall k\in\mathcal{K}$, are randomly drawn from the range of $[-\pi/4,\pi/4]$ and $[-\pi/6,\pi/6]$ with uniform distribution. $\kappa$ is set to 20dB to reflect the sparse nature of channels in typical large-scale antenna arrays. It can be observed that the conclusions drawn from small-scale systems as discussed previously, i.e., RSMA schemes outperform SDMA in both sum-rate and MMF, and SIC-free RSMA preserves most of the superiority of SIC-based RSMA, still holds in large-scale systems. Moreover, the proposed user grouping strategy generally enhances the performance of both SDMA and RSMA, with a more pronounced benefit for RSMA due to the high channel similarity within each user group.
@@ VII Conclusion
In this work, we aimed to investigate the fundamental limits of SIC-based and SIC-free RSMA under finite constellations. We first derived the rate expressions for SIC-based and SIC-free RSMA. We then proposed precoding optimization algorithms for the WSR and MMF optimization problems. To enable these optimization algorithms in large-scale systems with feasible computational burden, we introduced user grouping, null space projection and dimensionality reduction techniques. The numerical results show that RSMA schemes outperform SDMA under finite constellations and in both WSR or MMF evaluation, similar to the Gaussian distributed inputs. Furthermore, SIC-free RSMA preserves most of the superiority of SIC-based RSMA, which extends the conclusion in [19] from low-complexity precoders to fully optimized precoders, from sum-rate objective to WSR and MMF objectives, and from small-scale systems to large-scale systems.
@@ Appendix A Proof of Proposition 1
Suppose $\mathbf{P}$ is given and the private rates, $R_{\text{p},k},\forall k\in\mathcal{K}$, are not functions of $\mathbf{c}$, $\mathcal{P}_{1}$ is equivalent to
$\begin{split}\mathcal{P}_{8}:\;\;\underset{\mathbf{c}}{\text{max}}\>\>&\sum_{k\in\mathcal{K}}u_{k}C_{k}\\ \text{s.t.}\>\>\>\;&\sum_{k\in\mathcal{K}}C_{k}\leq R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K},\end{split}$ (43)
@@ Appendix B Proof of Proposition 2
$\mathcal{P}_{4}$ can be written as follows.
$\begin{split}\mathcal{P}_{9}:\;\;\max_{\mathbf{c}}&\;\;\min_{k}C_{k}+R_{\text{p},k}\\ \text{s.t.}&\;\;\sum_{k\in\mathcal{K}}C_{k}=R_{\text{c}}\\ &\;\;\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K}\\ \end{split}$ (44)
We introduce a slack variable, $\eta$, and convert $\mathcal{P}_{9}$ into
$\begin{split}\mathcal{P}_{10}:\;\;\max_{\mathbf{c},t}&\;\;\;\;\;\;\eta\\ \text{s.t.}&\;\;\sum_{k\in\mathcal{K}}C_{k}=R_{\text{c}}\\ &\;\;\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K}\\ &\;\;C_{k}+R_{\text{p},k}\geq\eta,\forall k.\end{split}$ (45)
Let $(\mathbf{c}^{\star},\eta^{\star})$ denote the solution of $\mathcal{P}_{10}$. Without loss of generality, we assume $R_{\text{p},1}<R_{\text{p},2}<...<R_{\text{p},K}$, and the solution of $\mathcal{P}_{8}$ leads to $C_{k}^{\star}>0$, $\forall k\in\{1,...,k^{\prime}\}$ with $1\leq k^{\prime}\leq K$, and $C_{k}^{\star}=0$, $\forall k\in\{k^{\prime}+1,...,K\}$. We have the following lemma.
@@
$C_{k}^{\star}+R_{\text{p},k}=\eta^{\star}$, $\forall k\in\{1,...,k^{\prime}\}$, and with some value for $\eta^{\star}$.
@@
It is trivial to see that $R_{\text{p},k}<\eta^{\star}$, $\forall k\in\{1,...,k^{\prime}\}$. Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be a partition of $\{1,...,k^{\prime}\}$. Suppose $(\mathbf{c}^{\star},\eta^{\star})$ leads to $C^{\star}_{k}+R_{\text{p},k}>\eta^{\star}$ for $k\in\mathcal{K}_{1}$, and $C^{\star}_{k}+R_{\text{p},k}=\eta^{\star}$ for $k\in\mathcal{K}_{2}$. Then, another choice of $\mathbf{c}$, denoted by $\mathbf{c}^{\diamond}$, can be found by $\mathbf{c}^{\diamond}=\alpha\mathbf{c^{\circ}}$, where $C^{\circ}_{k}=(C^{\star}_{k}-R_{\text{p},k}+\eta^{\star})/2$ for $k\in\mathcal{K}_{1}$, $C^{\circ}_{k}=C_{k}^{\star}$ for $k\in\mathcal{K}_{2}$, and $\alpha=R_{\text{c}}/\sum_{k\in\mathcal{K}}C^{\circ}_{k}$. It is trivial to see that $\alpha>1$. Therefore, $\min_{k\in\{1,...,k^{\prime}\}}C^{\circ}_{k}+R_{\text{p},k}>\min_{k\in\{1,...,k^{\prime}\}}C^{\star}_{k}+R_{\text{p},k}$, which indicates that $\mathbf{c}^{\star}$ is not the solution of $\mathcal{P}_{10}$. By contradiction, Lemma 2 is proved. ∎
@@
$\mathbf{c}^{\prime}+\mathbf{d}=\eta\mathbf{1}^{1\times k^{\prime}},$ (46)
where $\mathbf{d}=[R_{\text{p},1},...,R_{\text{p},k^{\prime}}]$, and $\mathbf{c}^{\prime}=[C_{1},...,C_{k^{\prime}}]$. Combining (46) and the constraint on $\sum_{k\in\mathcal{K}}C_{k}=R_{\text{c}}$, the global optimal solution of $\mathbf{c}^{\prime}$ and $\eta$, denoted by $\mathbf{c}^{\prime\star}$ and $\eta^{\star}$, can be then obtained by
$\mathbf{c}^{\prime\star}=\eta^{\star}\mathbf{1}^{1\times k^{\prime}}-\mathbf{d},$ (47)
@@
$\eta^{\star}=\frac{R_{\text{c}}+\mathbf{d}\mathbf{1}^{k^{\prime}\times 1}}{k^{\prime}}.$ (48)
Although $k^{\prime}$, i.e., the number of users whose are strictly served by the common stream, is still not known, we can test it using (47) and (48). This is because allocating the common stream rate based on (47) and (48) to more users than $k^{\prime}$ will lead to $\mathbf{c}^{\prime}$ containing negative entries, which is infeasible. Hence, by iteratively guessing the number of users to be served by the common stream, Algorithm 3 is obtained.
@@ Appendix C Transmission Mode Dictionaries Used For Numerical Results
The transmission mode dictionaries used to produce the numerical results in Section VI are depicted in Tables VI-VI.